A financial analyst has obtained the following:

Cash P200,000.00, Marketable securities P100,000.00, Accounts receivable, net P300,000.00, Inventories, net P480,000.00, Prepaid expenses P120,000.00, Accounts payable P250,000.00, Income Taxes P50,000.00, Accrued liabilities P100,000.00, Current portion of long term debt P200,000.00. Calculate the Company’s cash ratio
Answer #1

The formula of cash ratio is:

Cash ratio = Cash + Marketable securities / Current liabilities

Given: Cash = 200000, Marketable securities = 100000

Current liabilities = Accounts payable + Accrued liabilities + Income taxes payable + Current portion of long term debt (it is assumed here that income taxes given in the question is income tax payable)

Current liabilities = 250000 + 100000 + 50000 + 200000 = 500000

Now, putting these values in the cash ratio formula, we get,

Cash ratio = (200000 + 100000) / 500000

Cash ratio = 300000 / 500000 = 0.6